OpenStudy (anonymous):

.Write the equation of the transformed graph of sine with period pi that has been shifted horizontally to the right 3 units, has an amplitude of 3/4 , and has been flipped across the x-axis

OpenStudy (anonymous):

The period is normally 2 pi, which means if you want pi you need to multiply the inside by 2 (this divides the period in half). Then to shift to the right 3 units you need a -3 inside, and the amplitude is what you multiply it by so with a negative to flip it over the x axis: \[\sin(\psi) \rightarrow \frac{-3}{4} \sin(2(\psi-3))\]